# HYPE: Hyperbolic Entailment Filtering for Underspecified Images and Texts

### v1

@@ 1 Introduction
Figure 1: Example of HYPE filtering on the Datacomp small pool [1]. HYPE leverages both uni-modal specificity (text specificity $\epsilon_{t}$ and image specificity $\epsilon_{i}$) and cross-modal similarity (CLIP similarity $cos(\theta)$ as in this figure or negative Lorentizan distance $-d_{\mathcal{L}}$ can be used instead) to effectively identify and eliminate misalignment and underspecification issues on noisy image-text pairs. Figures (a-c) show instances where image-text pairs exhibit high alignment yet are flagged for exclusion due to insufficient specificity: (b) demonstrates low image specificity $\epsilon_{i}$, (c) illustrates low text specificity $\epsilon_{t}$, and (a) indicates low specificity in both aspects. Conversely, Figure (d) presents a scenario with high $\epsilon_{i}$ and $\epsilon_{t}$ but low $cos(\theta)$, highlighting a misalignment between the image and text, evidenced by the absence of an “elephant print”.
@@
The large datasets being created today, except private in-house datasets [7, 8, 9, 10], rely heavily on web-crawled documents by CommonCrawl [^1] https://commoncrawl.org/. As the scale of images and texts obtained from the web is tremendously large, each dataset employs different heuristics for reducing the size of the dataset. These heuristics include, for example, whether the text is a title from Wikipedia, whether it is written in English, and whether the image resolution is large enough [11, 12, 13, 14, 1]. Another rule-of-thumb is model-based filtering, usually based on the pre-trained CLIP [11] model, which determines if the given image and text are semantically aligned [13, 14, 1], or if the given image is similar to high-quality images from human-validated datasets, such as ImageNet [1].
Although CLIP-based filtering helps the verification of the semantic alignment between images and texts, we argue that only considering alignment is not enough criterion for high-quality data filtering. More specifically, we have to consider specificity of each data point. In this paper, we (informally) define alignment as whether a given image-text pair is sufficiently similar and specificity as whether a given unimodal data point contains sufficient information to be uniquely defined (i.e., specificity measures how each data point has semantically overlapping with other data points). A more formal definition will be described in Section 3.3. Figure 1 illustrates the concept of alignment and specificity. In the figure, we observe that the website screenshot and the URI are well-aligned, but the information of the screenshot and the URI are not sufficiently enough to be uniquely defined. Unfortunately, as CLIP-based filtering only considers alignment, it cannot filter out underspecified images and texts.
@@
    $\epsilon_{i}$, also can be used independently to induce a dataset to train image-only models. We show that the dataset filtered by $\epsilon_{i}$ trains a better image-only self-supervised model than the alignment-based filtering.
@@ 3.1 Overview of HYPE
Using the entailment cones, we define the “entailment loss” $\mathcal{L}_{e}(\mathbf{x},\mathbf{y})$ for the given image-text pair that measures whether the image (or text) is correctly placed on the entailment cone of the corresponding text (or image). Then, we define the “specificity” of each input by computing the average entailment loss on the dataset. Namely, the image specificity $\epsilon_{i}$ is defined as the average entailment loss, i.e., $\sum_{\mathbf{y}}\mathcal{L}_{e}(\mathbf{x},\mathbf{y})$, and the text specificity $\epsilon_{t}$ is defined similarly. $\epsilon_{i}$ and $\epsilon_{t}$ measure whether the given input is well described by the learned hyperbolic embedding space. We will provide a more rigorous mathematical definition in the latter section. Figure 3 shows examples of images and texts with low and high specificity values (i.e., $\epsilon_{i}$ and $\epsilon_{t}$, respectively). As shown in the figure, samples with smaller specificities are more generic and underspecified. For example, the low $\epsilon_{i}$ values of mobile phone or tower images denote their abundant potential relative captions in the dataset. Conversely, Dalmore whisky in the “Highest” category highlights the scarcity of descriptive texts without directly mentioning “Dalmore”, underscoring the metric’s effectiveness in distinguishing specificity. Similarly, the captions “pic” and “Picture” have low $\epsilon_{t}$ values as they are vague to describe a specific image.
@@ 3.2 Hyperbolic CLIP
Since $-d_{\mathcal{L}}$ can only be calculated on a manifold, not the tangent space, we need to map $\mathbf{v}$ in the tangent space to the manifold. Luckily, as MERU only deals with the tangent space of the origin $\mathbf{O}$, this exponential map can be simplified into:
$\mathbf{x}_{space}=\frac{sinh(\sqrt{c}\;\lVert\mathbf{v}_{space}\rVert)}{\sqrt{c}\;\lVert\mathbf{v}_{space}\rVert}\mathbf{v}_{space}$ (4)
@@ 3.3 Entailment Cone and Specificity
Now, we describe how we can measure specificity using hyperbolic embeddings. Note that $-d_{\mathcal{L}}$ also can perform as a filtering metric as a better alignment measure rather than the vanilla CLIP distance $cos(\theta)$. However, $-d_{\mathcal{L}}$ is also only able to measure alignment between images and texts but cannot measure how each image or text is specific. In this paper, we propose a new instance-wise filtering metric named specificity based on the concept of entailment. The concept of entailment has its roots in logic and linguistics, long before its incorporation into machine learning [54, 55]. In logic, entailment is a fundamental relationship where the truth of one statement guarantees the truth of another. In natural language processing, a number of tasks have been created to verify that the language model can properly capture this entailment relationship (i.e., semantic containment and exclusion): RTE [56, 57, 58, 59], MNLI [60], WNLI [61], etc., and these tasks form a significant part of the GLUE benchmark [29, 30]. Beyond natural language processing, tasks have also been created in the vision-and-language domain, such as SNLI-VE [62, 63], to evaluate cross-modal entailment relationships between images and text. The concept of an entailment cone emerges when we consider how entailment relationships can be represented in a vector space. The idea is that for a given concept or term represented by a vector, there exists a cone in the vector space within which all vectors that are semantically entailed by the original term fall.
@@ 3.4 Hyperbolic Entailment Filtering (HYPE)
Note that the original MERU by Desai et al. [15] was trained solely on the RedCaps [70] dataset. We added more clean data points to allow better filtering capability, as the findings of DFN [17] and our discussion in Section 4.1. We also note that the original MERU uses ViT-B/16 and ViT-L/16 backbones with their textual encoder having a hidden size of 512. Since DataComp [1] uses ViT-B/16 and ViT-L/14 for its baseline CLIP filtering method, we retrained MERU on ViT-B/16, which has a 512 textual encoder hidden size, and ViT-L/14, which has a 768 textual encoder hidden size, with the expanded dataset. The results of MERU re-training are shown in Table 2. Surprisingly, even when all the training hyperparameters, including the batch size, were the same as in the original MERU, and the training was done with fewer steps (ViT-B/16), the zero-shot performance of the reproduced MERU model was significantly better than that of the original MERU. All results in this paper are based on the hyperbolic embeddings obtained by our reproduced ViT-L/14 MERU.
We extract $\epsilon_{i}$, $\epsilon_{t}$, and $-d_{\mathcal{L}}$ for every sample in the target image-text dataset using our MERU model. For each sample, we also compute and store the ImageNet clustering-based image filter used by DataComp and the CLIP score $cos(\theta)$ of the ViT-L/14 CLIP. The ImageNet clustering-based image filter $c_{\text{IN}}$ is quantified as a value of 10 if included and 0 if not, enabling preferential use. Table 1 summarizes the statistics for the datasets tested in this paper. The $\text{HYPE}_{\text{score}}$ is obtained by linearly combining all the metrics with equal weight as defined in Eqn. 1.
@@ 4.2 Image-Text Contrastive Pre-training
Second, we observe that HYPE can be combined with the other filtering methods. As our specificity metric is single-modality filtering and orthogonal to cross-modality filtering, such as CLIP score filtering, all other baselines rely on, it can properly filter underspecified examples as shown in Figure 1. This characteristic helps us to mark the first rank in the DataComp small and medium track by combining HYPE with DFN.
@@ 4.3 Image-Only Contrastive Pre-training
Table 5: ImageNet linear probing classification accuracies of ViT-S. The table compares $cos(\theta)$ and $\epsilon_{i}$ on inducing various size image-only datasets for image-only self-supervised learning methods: SimCLR [79] and MoCo-v3 [80].
Model Filtering Metric Dataset Size
0.13M 0.32M 0.64M 1.28M
SimCLR [79] $cos(\theta)$ 47.06 45.47 52.31 49.68
$\epsilon_{i}$ 53.30 50.89 57.49 54.55
MoCo-v3 [80] $cos(\theta)$ 39.00 45.00 51.10 53.40
$\epsilon_{i}$ 44.70 51.60 56.80 59.70
@@ Appendix A Histograms
Figure A.1: In this paper, we examine four key metrics: text specificity ($\epsilon_{t}$), image specificity ($\epsilon_{i}$), negative Lorentzian distance ($-d_{\mathcal{L}}$), and CLIP cosine similarity ($cos(\theta)$). For each metric, we present histograms to illustrate their distribution across various subsets of the Datacomp medium data pool. These subsets are color-coded for clarity: DFN [17] is shown in blue, CIDS [73] in orange, our method (HYPE) in green, and the no-filter condition in red. Additionally, we highlight their average values for each metric with vertical dotted lines in the respective histograms. Each method results in a different amount of data in the subset, so for ease of comparison, the y-axis shows the relative percentage rather than the count.
In Figure A.1, we employed histograms to visually examine the alignment of each filtering method’s subset with the metrics investigated in our study: text specificity ($\epsilon_{t}$), image specificity ($\epsilon_{i}$), negative Lorentzian distance ($-d_{\mathcal{L}}$), and CLIP cosine similarity ($cos(\theta)$). For all metrics except CLIP cosine similarity, a distinct hierarchy emerged among the methods: Unfiltered was the least aligned, followed by CIDS, then DFN, and HYPE showing the highest alignment. The histogram results closely reflect the uniform sampling approach adopted for the subsets, as detailed in the uniform column of Table 4. Uniform sampling ensures each sample in the subset is selected with equal probability. However, when the sampling method strays from this uniform approach, such as the duplication sampling in the CIDS [73], the histograms will lean towards data points that are sampled more frequently. Such deviations could significantly impact the representativeness of the histograms.
@@
Appendix B More Qualitative Results
@@
In the remaining sections of the supplementary material, additional examples corresponding to the specificities highlighted in Figure 3 are presented. To enhance understanding of the relationship between $cos(\theta)$, $\epsilon_{t}$, and $\epsilon_{i}$, as initially introduced in Figure 1, we arranged the samples to be ordered from the left to the right, progressing from the lowest to the highest $cos(\theta)$ values. This layout aids in visualizing the concept mentioned in Figure 1, demonstrating that with increasing $cos(\theta)$ values, image samples tend to contain texts, aligning more closely with those used in the OCR tasks.
@@
Figure B.1: More example texts from Datacomp small that fall within the percentile of $p(\epsilon_{t})<1\%$. To ensure uniqueness, we removed duplicates, so each text, despite appearing as empty space in the figures, contains distinct characters. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.2: More example texts from Datacomp small that fall within the percentile of $19\%<p(\epsilon_{t})<21\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.3: More example texts from Datacomp small that fall within the percentile of $39\%<p(\epsilon_{t})<41\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.4: More example texts from Datacomp small that fall within the percentile of $59\%<p(\epsilon_{t})<61\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.5: More example texts from Datacomp small that fall within the percentile of $79\%<p(\epsilon_{t})<81\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.6: More example texts from Datacomp small that fall within the percentile of $99\%<p(\epsilon_{t})$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.7: More example images from Datacomp small that fall within the percentile of $p(\epsilon_{i})<1\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.8: More example images from Datacomp small that fall within the percentile of $19\%<p(\epsilon_{i})<21\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.9: More example images from Datacomp small that fall within the percentile of $39\%<p(\epsilon_{i})<41\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.10: More example images from Datacomp small that fall within the percentile of $59\%<p(\epsilon_{i})<61\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.11: More example images from Datacomp small that fall within the percentile of $79\%<p(\epsilon_{i})<81\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure B.12: More example images from Datacomp small that fall within the percentile of $99\%<p(\epsilon_{i})$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.

### v2

@@ 1 Introduction
Figure 1: Example of HYPE filtering on the Datacomp small pool [20]. HYPE leverages both uni-modal specificity (text specificity $\epsilon_{t}$ and image specificity $\epsilon_{i}$) and cross-modal similarity (CLIP similarity $cos(\theta)$ as in this figure or negative Lorentzian distance $-d_{\mathcal{L}}$ can be used instead) to effectively identify and eliminate misalignment and underspecification issues on noisy image-text pairs. Figures (a-c) show instances where image-text pairs exhibit high alignment yet are flagged for exclusion due to insufficient specificity: (b) demonstrates low image specificity $\epsilon_{i}$, (c) illustrates low text specificity $\epsilon_{t}$, and (a) indicates low specificity in both aspects. Conversely, Figure (d) presents a scenario with high $\epsilon_{i}$ and $\epsilon_{t}$ but low $cos(\theta)$, highlighting a misalignment between the image and text, evidenced by the absence of an “elephant print”.
@@
The large datasets being created today, except private in-house datasets [62, 79, 74, 54], rely heavily on web-crawled documents by CommonCrawl. As the scale of images and texts obtained from the web is tremendously large, each dataset employs different heuristics for reducing the size of the dataset. These heuristics include, for example, whether the text is a title from Wikipedia, whether it is written in English, and whether the image resolution is large enough [53, 73, 58, 57, 20]. Another rule-of-thumb is model-based filtering, usually based on the pre-trained CLIP [53] model, which determines if the given image and text are semantically aligned [58, 57, 20], or if the given image is similar to high-quality images from human-validated datasets, such as ImageNet [20].
Although CLIP-based filtering helps verify the semantic alignment between images and texts, we argue that alignment alone is insufficient for high-quality data filtering. More specifically, we should consider specificity of each data point. Here, we (informally) define alignment as whether a given image-text pair is sufficiently similar and specificity as whether a given unimodal data point contains sufficient information to be uniquely defined (i.e., specificity measures how each data point has semantically overlapping with other data points). A more formal definition will be described in Sec. 3.3. Figure 1 illustrates the concept of alignment and specificity. In the figure, the website screenshot and the URI are well-aligned, but the information of the screenshot and the URI are not sufficiently enough to be uniquely defined. Unfortunately, as CLIP-based filtering only considers alignment, it cannot filter out underspecified images and texts.
@@
    $\epsilon_{i}$ can be independently applied to induce a dataset for training image-only self-supervised models, showing superior performance compared to alignment-based filtering.
@@ 3.1 Overview of HYPE
Using the entailment cones, we define the “entailment loss” $\mathcal{L}_{e}(\mathbf{x},\mathbf{y})$ for the given image-text pair that measures whether the image $\mathbf{y}$ is correctly placed on the entailment cone of the corresponding text $\mathbf{x}$. Then, we define the “specificity” of each input by computing the average entailment loss on the dataset. The image specificity $\epsilon_{i}$ is defined as the average entailment loss, i.e., $\sum_{\mathbf{x}}\frac{\mathcal{L}_{e}(\mathbf{x},\mathbf{y})}{M}$, and the text specificity $\epsilon_{t}$ is defined similarly, $\sum_{\mathbf{y}}\frac{\mathcal{L}_{e}(\mathbf{x},\mathbf{y})}{M}$. $\epsilon_{i}$ and $\epsilon_{t}$ measure whether the learned hyperbolic embedding space describes the given input well. We will provide a more rigorous mathematical definition in the latter section. Figure 3 shows examples of images and texts with low and high specificity values (i.e., $\epsilon_{i}$ and $\epsilon_{t}$, respectively). As shown in the figure, samples with smaller specificities are more generic and underspecified. For example, the low $\epsilon_{i}$ values of mobile phone or tower images denote their abundant potential relative captions in the dataset. Conversely, Dalmore whisky in the “Highest” category highlights the scarcity of descriptive texts without directly mentioning “Dalmore”, underscoring the metric’s effectiveness in distinguishing specificity. Similarly, the captions “pic” and “Picture” have low $\epsilon_{t}$ values as they are vague to describe a specific image.
@@ 3.2 Hyperbolic CLIP
As $-d_{\mathcal{L}}$ can only be calculated on a manifold, not the tangent space, we need to map $\mathbf{v}$ in the tangent space to the manifold. Luckily, as MERU only deals with the tangent space of the origin $\mathbf{O}$, this exponential map can be simplified into:
$\mathbf{x}_{space}=\frac{\sinh(\sqrt{c}\;\lVert\mathbf{v}_{space}\rVert)}{\sqrt{c}\;\lVert\mathbf{v}_{space}\rVert}\mathbf{v}_{space}$ (4)
@@ 3.3 Entailment Cone and Specificity
Now, we describe how we can measure specificity using hyperbolic embeddings. Note that $-d_{\mathcal{L}}$ also can perform as a filtering metric as a better alignment measure rather than the vanilla CLIP distance $cos(\theta)$. However, $-d_{\mathcal{L}}$ can only measure alignment between images and texts but cannot measure how each image or text is specific. This paper proposes a new instance-wise filtering metric named specificity based on the concept of entailment. The concept of entailment has its roots in logic and linguistics, long before its incorporation into machine learning [66, 64]. In logic, entailment is a fundamental relationship where the truth of one statement guarantees the truth of another. In natural language processing (NLP), a number of tasks have been created to verify that the language model can properly capture this entailment relationship (i.e., semantic containment and exclusion): RTE [14, 3, 24, 5], MNLI [70], WNLI [41], etc., and these tasks form a significant part of the GLUE benchmark [68, 69]. Beyond NLP, tasks have also been created in the vision-and-language domain, such as SNLI-VE [72, 71], to evaluate cross-modal entailment relationships between images and text. The concept of an entailment cone emerges when we consider how entailment relationships can be represented in a vector space. The idea is that for a given concept represented by a vector, there exists a cone in the vector space within which all vectors that are semantically entailed by the original term fall.
@@ 3.4 Hyperbolic Entailment Filtering (HYPE)
Note that the original MERU by Desai et al. [16] was trained solely on the RedCaps [15] dataset. We added more clean data points to allow better filtering capability, as the findings of DFN [19] and our discussion in Sec. 4.1. We also note that the original MERU uses ViT-B/16 and ViT-L/16 backbones with their textual encoder having a hidden size of 512. Since DataComp [20] uses ViT-B/16 and ViT-L/14 for its baseline CLIP filtering method, we retrained MERU on ViT-B/16, which has a 512 textual encoder hidden size, and ViT-L/14, which has a 768 textual encoder hidden size, with the expanded dataset. The results of MERU re-training are shown in Table 2. Surprisingly, even when all the training hyperparameters, including the batch size, were the same as in the original MERU, and the training was done with fewer steps (ViT-B/16), the zero-shot performance of the reproduced MERU model was significantly better than that of the original MERU. All results in this paper are based on the hyperbolic embeddings obtained by our reproduced ViT-L/14 MERU.
We extract $\epsilon_{i}$, $\epsilon_{t}$, and $-d_{\mathcal{L}}$ for every sample in the target image-text dataset using our MERU model. For each sample, we also compute and store the ImageNet clustering-based image filter used by DataComp and the CLIP score $cos(\theta)$ of the ViT-L/14 CLIP. The clustering-based filter $c_{\text{IN}}$ is quantified as a value of 10 if included and 0 if not, enabling preferential use. Table 1 summarizes the statistics for the datasets tested in this paper. The $\text{HYPE}_{\text{score}}$ is obtained by linearly combining all the metrics with equal weight as defined in Eqn. 1.
@@ 4.2 Image-Text Contrastive Pre-training
Second, HYPE can be combined with the other methods. As our specificity metric is single-modality filtering and orthogonal to cross-modality filtering, such as CLIP filtering, all other baselines rely on, it can properly filter underspecified examples as shown in Figure 1. This characteristic helps us to mark the first rank in the DataComp small and medium track by combining HYPE with DFN.
Figure 6: Filtering network IN-ZS acc. vs. Induced networks IN-ZS acc. (overlaid on the figure of the DFN paper [19])
Additionally, we trained two more B/16 models with different training seen samples, 128M and 256M, whose IN-ZS accuracies are 42.3% and 47.6%. With these models, we report their correlation to DataComp medium’s IN-ZS accuracy as described in [19]. As shown in Figure 6, a better-performing MERU model consistently induces a more effective filtering network, evidenced by improved zero-shot accuracy. This upward trend suggests that further improvements in MERU could lead to even more effective filtering, which is not observed in the downward trend of Euclidean CLIP.
@@ 4.3 Image-Only Contrastive Pre-training
Table 5: ImageNet-1K linear probing classification accuracies of ViT-S. The table compares $cos(\theta)$ and $\epsilon_{i}$ on inducing various size image-only datasets (DataComp Medium) for image-only self-supervised learning methods: SimCLR [9] and MoCo-v3 [10].
Model Filtering Metric Dataset Size 0.13M 0.32M 0.64M 1.28M SimCLR [9] $cos(\theta)$ 47.06 45.47 52.31 49.68 $\epsilon_{i}$ 53.30 50.89 57.49 54.55 MoCo-v3 [10] $cos(\theta)$ 39.00 45.00 51.10 53.40 $\epsilon_{i}$ 44.70 51.60 56.80 59.70
@@ Appendix 0.A Histograms
Figure 0.A.1: In this paper, we examine four key metrics: text specificity ($\epsilon_{t}$), image specificity ($\epsilon_{i}$), negative Lorentzian distance ($-d_{\mathcal{L}}$), and CLIP cosine similarity ($cos(\theta)$). For each metric, we present histograms to illustrate their distribution across various subsets of the Datacomp medium data pool. These subsets are color-coded for clarity: DFN [19] is shown in blue, CIDS [76] in orange, our method (HYPE) in green, and the no-filter condition in red. Additionally, we highlight their average values for each metric with vertical dotted lines in the respective histograms. Each method results in a different amount of data in the subset, so for ease of comparison, the y-axis shows the relative percentage rather than the count.
In Figure 0.A.1, we employed histograms to visually examine the alignment of each filtering method’s subset with the metrics investigated in our study: text specificity ($\epsilon_{t}$), image specificity ($\epsilon_{i}$), negative Lorentzian distance ($-d_{\mathcal{L}}$), and CLIP cosine similarity ($cos(\theta)$). For all metrics except CLIP cosine similarity, a distinct hierarchy emerged among the methods: Unfiltered was the least aligned, followed by CIDS, then DFN, and HYPE showing the highest alignment. The histogram results closely reflect the uniform sampling approach adopted for the subsets, as detailed in the uniform column. Uniform sampling ensures each sample in the subset is selected with equal probability. However, when the sampling method strays from this uniform approach, such as the duplication sampling in the CIDS [76], the histograms will lean towards data points that are sampled more frequently. Such deviations could significantly impact the representativeness of the histograms.
@@
Appendix 0.B More Qualitative Results
@@
In the remaining sections of the supplementary material, additional examples corresponding to the specificities are presented. To enhance understanding of the relationship between $cos(\theta)$, $\epsilon_{t}$, and $\epsilon_{i}$, we arranged the samples to be ordered from the left to the right, progressing from the lowest to the highest $cos(\theta)$ values. This layout aids in visualizing the concept mentioned in the paper, demonstrating that with increasing $cos(\theta)$ values, image samples tend to contain texts, aligning more closely with those used in the OCR tasks.
@@
Figure 0.B.1: More example texts from Datacomp small that fall within the percentile of $p(\epsilon_{t})<1\%$. To ensure uniqueness, we removed duplicates, so each text, despite appearing as empty space in the figures, contains distinct characters. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.2: More example texts from Datacomp small that fall within the percentile of $19\%<p(\epsilon_{t})<21\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.3: More example texts from Datacomp small that fall within the percentile of $39\%<p(\epsilon_{t})<41\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.4: More example texts from Datacomp small that fall within the percentile of $59\%<p(\epsilon_{t})<61\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.5: More example texts from Datacomp small that fall within the percentile of $79\%<p(\epsilon_{t})<81\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.6: More example texts from Datacomp small that fall within the percentile of $99\%<p(\epsilon_{t})$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.7: More example images from Datacomp small that fall within the percentile of $p(\epsilon_{i})<1\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.8: More example images from Datacomp small that fall within the percentile of $19\%<p(\epsilon_{i})<21\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.9: More example images from Datacomp small that fall within the percentile of $39\%<p(\epsilon_{i})<41\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.10: More example images from Datacomp small that fall within the percentile of $59\%<p(\epsilon_{i})<61\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.11: More example images from Datacomp small that fall within the percentile of $79\%<p(\epsilon_{i})<81\%$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.
@@
Figure 0.B.12: More example images from Datacomp small that fall within the percentile of $99\%<p(\epsilon_{i})$. The examples are sequentially sampled in ascending order of CLIP scores within the given percentile pool.